OpenStudy (anonymous):

Part A: Divide (10x4y3 + 5x3y2 - 15x2y - 25x2y4) by -5x2y. Show your work, and justify each step. (6 points) Part B: How would your answer in Part A be affected if the x2 variable in the denominator was just an x? (2 points) Part C: What is the degree and classification of the polynomial you got in Part A? (2 points)

OpenStudy (anonymous):

helllllpppppppppppp

OpenStudy (anonymous):

\(10x4y^3 + 5x^3y^2 - 15x2y - 25x2y^4\)-> factor out the common terms first.

OpenStudy (abmon98):

\[10x^4y^3/-5x^2y=-2x^2y^2\] \[5x^3y^2/-5x^2y=-xy\] \[-15x^2y/-5x^2y=3\] \[-25x^2y^4/-5x^2y=5y^3\] \[-2x^2y^2-xy+3+5y^3 \]

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

for A